Problem 5 Accurately sketch the following function and evaluate the integral 15 Points - (s(6-(4) W 3W x(t) A tri 3W W A tri​Answer both problems please.
Answer #1


Crivera thnat -30 Li A tilt 8 (t-toylt It tto) Atrictls) Bb x1t2= Atri/w/4 2)8 - しt)こ Ari (\1u)81tw/u)+Ati (3/u) 81t-3w/4) trilt) w- ぐつ1-も 1 tnciu= 1-1 3/4 1-3/12Y ta(31u) lt)- 3A 81Le anpio taat SI-do) Ststf0) 1 2 8-S0)-S Ifo) 2 simersot (costb02)+Sirrj cosCbi-0